The Ratio of the side lengths of a triangle is 4:7:9. The perimeter of the triangle is 120 feet. Find the side lengths. Then classify the triangle by its side lengths.
Answer by stanbon(75887)

The Ratio of the side lengths of a triangle is 4:7:9. The perimeter of the triangle is 120 feet. Find the side lengths. Then classify the triangle by its side lengths.
---
Equation:
4x + 7x + 9x = 120
20x = 120
x = 6
---
Lengths: 4*6 = 24 ; 7*6 = 42 ; 9*6 = 54
-----
Use the Law of Cosines to find angle sizes so you can classify the triangle.
e.g. Cos(A) = [24^2+42^2-54^2]/(2*24*42) = -0.2857
A = cos^-1(-0.285) = 106 degrees
etc.
-------------------
Cheers,
Stan H.